A hydrogen atom, initially in the ground state, ab
Answer #1

we have the formula of wave number

wave number = 1/\lambda = 109677((1/n1)^2 - (1/n2)^2)

1/97.5 * 10^-7 cm = 109677((1/1^2) - (1/n2)^2)

n2 = 3.93 which is nearly equal to 4